Jodie Foster Care Homes Inc. shows the following data Total Stockholders Year Net Income Assets 20x1 S 138,000 S 2,280,000 5 959,000 20X2 20x3 20x4 Equity Total Debt 53,000 185,000 2,440,000 1,520,000 228,000 2,880,000 1,950,000 1321,000 996,000 920,000 730,000 2,110,000 1,114,000 a-1. Compute the ratio of net income to total assets for each year. (Input your answers as a percent rounded to 2 decimal places.) Year Return on 20x1 20X2 20x3 20X4 Assets a-2. What is the trend in the net income to total assets ratio? O Strong downward movement Strong upward movement b-1. Compute the ratio of net income to stockholders equity for each year. (Input your answers as a percent rounded to 2 decimal places.) Return on Year Stockholders 20x1 20X2 20X3 20X4 b-2. What is the trend in the net income to stockholders equity ratio? O Strong downward movement Strong upward movement
Answer #1
Answer = 1 A)
STEP 1: FORMULA
Net income to total assets = Net Income / Total Assets
STEP 2: CALCULATION OF THE RATIO
2001 2002 2003 2004
Net income to total assets =
Net Income = $          1,38,000 $        1,53,000 $               1,85,000 $       2,28,000
Divide By "/" By
Total Assets $        22,80,000 $     21,10,000 $             24,40,000 $    26,80,000
Net income to total assets =                       0.06                    0.07                            0.08                   0.09
Net income to total assets in % 6.05% 7.25% 7.58% 8.51%
STEP 3: ANSWER
Net income to total assets =
20X1 6.05%
20X2 7.25%
20X3 7.58%
20X4 8.51%
Answer = A -2)
Answer = Option 2 = Strong Upward Movement
Answer = B-1)
Return on Net income to ShareholdersEquity = Net Income / Shareholder's Equity
CALCULATION OF THE RATIO
2001 2002 2003 2004
Return on Net income to ShareholdersEquity =
Net Income = $          1,38,000 $        1,53,000 $               1,85,000 $       2,28,000
Divide By "/" By
Total Assets $          9,59,000 $     11,14,000 $             15,20,000 $    19,50,000
Return on Net income to ShareholdersEquity =                       0.14                    0.14                            0.12                   0.12
Return on Net income to ShareholdersEquity % = 14.39% 13.73% 12.17% 11.69%
STEP 3: ANSWER
Return on Net income to ShareholdersEquity =
20X1 14.39%
20X2 13.73%
20X3 12.17%
20X4 11.69%
Answer = B-2)
Answer = Option 1 = Strong Downward Movement